# PID Accelerated Temporal Difference Algorithms

### v1

@@ 5 Gain Adaptation
The proper choice of controller gains is critical to both convergence and acceleration of the proposed algorithms. While it is possible to treat the gains as hyperparameters and tune them like any other hyperparameter, we address this by designing an automatic gain adaptation algorithm that tunes them on the fly during the runtime of the algorithm.
@@
To avoid the difficulty of the inner product term, we modify the update rule of the gains at iteration $t$ to minimize $(\text{BR}^{\pi}V_{t+1})(X_{t})^{2}$ instead of $\left\lVert\text{BR}^{\pi}V_{t+1}\right\rVert_{2}^{2}$. This modification is similar to performing stochastic gradient descent instead of gradient descent. Instead of defining the loss over the whole state space, we consider the loss on a single sampled state. Consequently, the update depends only on the state and action of the current sample $(X_{t},A_{t},R_{t},X_{t}^{\prime})$. We get the following update:
@@
When calculating $\frac{\partial V_{t+1}(X_{t})}{\partial\kappa_{\cdot}}$, we further ignore the effect of gains on $V_{t}$, setting $\frac{\partial V_{t}}{\partial\kappa_{\cdot}}\approx 0$, and also drop the learning rate $\mu(N_{t}(X_{t}))$ to absorb it into $\eta$. These derivatives can be simply calculated based on (3) and are given in Appendix C. Finally, the normalization term is estimated by keeping an exponential moving average with smoothing factor $\lambda$ of the square of estimates of $\left\lVert\text{BR}^{\pi}V_{t}\right\rVert_{2}$ in past iterations. The final algorithm is shown in Algorithms 1 and 2 in Appendix C.
@@ 6 Empirical Results
Figure 1: Comparison of PID TD Learning with Conventional TD Learning in Chain Walk (left) and Cliff Walk (right) with $\gamma=0.9$. Each curve is average over 80 runs. Shaded areas show the standard error.
We empirically compare PID TD Learning and PID Q-Learning with their conventional counterparts. We conduct experiments in the 50-state Chain Walk environment with 2 actions (Farahmand and Ghavamzadeh, 2021), the Cliff Walk environment with $6\times 6$ states and 4 actions (Rakhsha et al., 2022), and randomly generated Garnet MDPs with 50 states and 3 actions (Bhatnagar et al., 2009). Detailed descriptions of these environments and the policies evaluated can be found in Appendix D. Samples are taken i.i.d from the environments. We measure the error of value functions $V_{t}$ and $Q_{t}$ for PE and Control problems by their normalized error defined as $\left\lVert V_{t}-V^{\pi}\right\rVert_{1}/\left\lVert V^{\pi}\right\rVert_{1}$ and $\left\lVert Q_{t}-Q^{*}\right\rVert_{F}/\left\lVert Q^{*}\right\rVert_{F}$, respectively.
@@
In Figure 1 we compare PID TD Learning with TD Learning when the gains are fixed and $\gamma=0.9$. In this case the acceleration depends on the choice of gains and the environment. We observe that we can achieve a drastic acceleration Cliff Walk, and a minor acceleration in Chain Walk. We further investigate the speed-up achieved by PID TD Learning in Cliff Walk. In Figure 2, we observe that with Gain Adaptation and $\gamma=0.999$, we achieve a significant acceleration without the need to tune the controller gains. Figure 2 also shows how Gain Adaptation has increased the gains from their initial values, which was shown to be useful in Figure 1.
@@
Figure 2: PID TD Learning with Gain Adaptation in Cliff Walk with $\gamma=0.999$. (Left) Comparison of value errors of PID TD Learning with TD Learning. Each curve is average over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 3: PID Q-Learning with Gain Adaptation in Chain Walk with $\gamma=0.999$. (Left) Comparison of value errors of PID Q-Learning with Q-Learning. Each curve is average over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 4: Comparison of PID Accelerated algorithms with the conventional ones for PE (Left) and Control (Right) problems in randomly generated Garnet environments with $\gamma=0.99$. Each curve is average of 80 MDPs. Shaded area shows standard error.
@@ 7 Related Work
The I component in our PID controller has a high-level similarity to the Momentum VI introduced by Vieillard et al. (2020). Instead of using the I component to update the values, however, they use a smoothed version of the value function to create the policy from which actions are sampled from. Their approach is limited to the Control setting and doesn’t address the bootstrapping error.
Gain adaptation in general has a long history in the RL and closely-related literature. Kesten (1958) used an adaptive mechanism in the context of stochastic approximation in the 1950s. They describe a method for choosing the learning rate of SA that is very similar to the P component of the gain adaptation procedure we naturally derive. However, the algorithm is ad hoc in nature, and is not compatible with function approximation in any natural way.
First order methods of adapting hyperparameters have been proposed, including IDBD (Sutton, 1992), the recent RL focused variant TIDBD (Kearney et al., 2018), and SMD (Schraudolph, 1999) which all tune learning rates by finding the gradient with respect to the history of errors. We refer to Sutton (2022) for a more in-depth history and overview of such techniques. These approaches are limited to controlling only the learning rate of the procedure and thus only attacking the error from sampling, not the bootstrapping error.
Aside from PID VI, there is a growing literature on accelerated dynamic programming methods in RL, including Accelerated VI (Goyal and Grand-Clément, 2023), Anderson VI (Geist and Scherrer, 2018), Anchored VI (Lee and Ryu, 2023), and the model-based approaches of OSVI (Rakhsha et al., 2022) and MoCoVI (Rakhsha et al., 2024). Other recent attempts at accelerating RL at the level of Bellman updates that take different perspectives include Speedy Q-Learning (Ghavamzadeh et al., 2011), a variant of Q-Learning with better PAC guarantees, and Zap Q-Learning (Devraj and Meyn, 2017) which is provably more stable, as well as a Zap Q-Learning variant for non-linear function approximation (Chen et al., 2020a). Zap Q-Learning requires significantly more computation than PID Q-Learning, being a second-order optimization method with a matrix inversion at each step.
@@ 8 Conclusion
We showed how the recent advances in accelerated planning and dynamic programming, specifically the PID Value Iteration algorithm, can be used to design algorithms for the RL setting. The proposed PID Temporal Difference Learning and PID Q Learning algorithms are accompanied with a gain adaptation mechanism, which tunes their hyperparameters on the fly. We provided theoretical analysis as well as empirical studies of these algorithms.
One limitation of the current work is that the proposed algorithms are only developed for finite MDPs where the value function, and all relevant quantities, can be represented exactly. For large MDPs, for example with continuous state spaces, we need to use function approximation. Developing PID TD and PID Q Learning with function approximation is therefore one important future direction.
Another limitation of this work is that the gain adaptation procedure, even though empirically reliable, does not come with a convergence guarantee. Another interesting research direction is then to develop gain adaptation procedure with convergence guarantee.
Finally, this work shows that the dynamics of RL can be significantly influenced by the PID controller, one of the simplest controller in the arsenal of control engineering. Developing Planning and RL algorithms based on more sophisticated controllers is another promising research direction.
@@
In this section, we present the proof of Theorem 1. We first present some notations, and then the assumptions on the learning rate schedule $\mu$ and sequence of visited samples $(X_{t})_{t\geq 0}$. Let $r^{\pi}\in\mathbb{R}^{n}$ be the vector of the expected immediate rewards of following the policy at each state.
@@
Examples of learning rate schedules that satisfy Assumption 1 includes $\mu(t)=\frac{1}{t+1}$. The next assumption is on balanced update of states.
@@
We show the claim by applying the result by Borkar and Meyn (2000, Theorem 2.5) to our algorithm. One can verify that the assumptions on the learning rate, update sequence, and noise are satisfied due to Assumption 1,2 and Lemma 1.
Borkar and Meyn (2000) require the origin point to be an asymptotically stable equilibrium of ODE
$\dot{u}(t)=(A_{g}^{\pi}-I)u(t),$
which is satisfied due to the assumption on the eigenvalues of $A_{g}^{\pi}-I$ and the fact that the solution of above ODE is $\exp{[(A_{g}^{\pi}-I)t]}u_{0}$ (Teschl, 2012).
@@
$\dot{u}(t)=L_{g}^{\pi}u(t)-u(t)=(A_{g}^{\pi}-I)u(t)+b_{g}^{\pi}.$
is $-(A_{g}^{\pi}-I)^{-1}b_{g}^{\pi}$ which is equal to $\tilde{V}^{\pi}$ due to $\tilde{V}^{\pi}=L_{g}^{\pi}\tilde{V}^{\pi}=A_{g}^{\pi}\tilde{V}^{\pi}+b_{g}^{\pi}$. Now we conclude that PID TD Learning converges to $\tilde{V}^{\pi}$.
@@
The existence of $S^{\prime}$ such that $\left\lVert A\right\rVert_{2,S^{\prime}}\leq\rho(A)+\delta$ is a consequence of the proof of Theorem 4.4 in (Householder, 1958). Due to equivalence of norms in finite dimensions, there exists $u>0$ such that for any $v\in\mathbb{R}^{d}$, $\left\lVert v\right\rVert_{2,S^{\prime}}\leq u\left\lVert v\right\rVert_{\infty}$. Define $S=\frac{1}{u}S^{\prime}$. Consequently, $\left\lVert v\right\rVert_{2,S}\leq\left\lVert v\right\rVert_{\infty}$ for any $v$. Now from Theorem 2.10 in (Householder, 1958), we have $\left\lVert A\right\rVert_{2,S^{\prime}}=\left\lVert S^{\prime}AS^{\prime-1}\right\rVert_{2}$ and $\left\lVert A\right\rVert_{2,S}=\left\lVert SAS^{-1}\right\rVert_{2}$ which means
@@
Assume a dataset $\{(x,R_{x},X^{\prime}_{x})\}_{x}$ is given which for each state $x$ contains the random reward $R_{x}\sim\mathcal{R}^{\pi}(x)$ and $X^{\prime}_{x}\sim{\mathcal{P}}^{\pi}(\cdot|x)$. For an arbitrary value function $V$, define $W\colon{\mathcal{X}}\to\mathbb{R}$ as
$W(x)\triangleq R_{x}+\gamma V(X^{\prime}_{x})-r^{\pi}(x)-\gamma{\mathcal{P}}^{\pi}V.$
@@
The result follows from the fact that for any random vector $Z\in\mathbb{R}^{m}$,
$\mathbb{E}\left[\left\lVert Z\right\rVert_{\infty}^{2}\right]\leq\mathbb{E}\left[\sum_{x}Z(x)^{2}\right]=\sum_{x}E\left[Z(x)^{2}\right].$

### v2

@@ 5 Gain Adaptation
The proper choice of controller gains is critical to both convergence and acceleration of our proposed algorithms. While it is possible to treat the gains as hyperparameters and tune them like any other hyperparameter, we address this by designing an automatic gain adaptation algorithm that tunes them on the fly during the runtime of the algorithm.
@@
To avoid the difficulty of the inner product term, we modify the update rule of the gains at iteration $t$ to minimize $(\text{BR}^{\pi}V_{t+1})(X_{t})^{2}$ instead of $\left\lVert\text{BR}^{\pi}V_{t+1}\right\rVert_{2}^{2}$. This modification is similar to performing stochastic gradient descent instead of gradient descent. Instead of defining the loss over the whole state space, we consider the loss on a single sampled state. Consequently, the new term only depends on the values at $X_{t}$. We get the following update:
@@
When calculating $\frac{\partial V_{t+1}(X_{t})}{\partial\kappa_{\cdot}}$, we further ignore the effect of gains on $V_{t}$, setting $\frac{\partial V_{t}}{\partial\kappa_{\cdot}}\approx 0$, and also drop the learning rate $\mu(N_{t}(X_{t}))$ to absorb it into $\eta$. These derivatives can be calculated based on (3) and are given in Appendix C. Finally, the normalization term is estimated by keeping an exponential moving average with smoothing factor $\lambda$ of the square of estimates of the Bellman Residual in the past iterations. The detailed version of PID TD Learning and PID Q-Learning with gain adaptaion is shown in Algorithms 1 and 2 in Appendix C.
@@ 6 Empirical Results
Figure 1: Comparison of PID TD Learning with Conventional TD Learning in Chain Walk (left) and Cliff Walk (right) with $\gamma=0.99$. Each curve is averaged over 80 runs. Shaded areas show the standard error.
We empirically compare PID TD Learning and PID Q-Learning with their conventional counterparts. We conduct experiments in the 50-state Chain Walk environment with 2 actions (Farahmand and Ghavamzadeh, 2021), the Cliff Walk environment with $6\times 6$ states and 4 actions (Rakhsha et al., 2022), and randomly generated Garnet MDPs with 50 states and 3 actions (Bhatnagar et al., 2009). Detailed descriptions of these environments and the policies evaluated can be found in Appendix D. For each sample $(X_{t},A_{t},R_{t},X_{t}^{\prime})$, we choose $X_{t}$ uniformly at random and $A_{t}$ is chosen according to $\pi$ (for PE) or at random (for Control). We measure the error of value functions $V_{t}$ and $Q_{t}$ for PE and Control problems by their normalized error defined as $\left\lVert V_{t}-V^{\pi}\right\rVert_{1}/\left\lVert V^{\pi}\right\rVert_{1}$ and $\left\lVert Q_{t}-Q^{*}\right\rVert_{F}/\left\lVert Q^{*}\right\rVert_{F}$, respectively, where $\left\lVert Q\right\rVert_{F}\triangleq(\sum_{x,a}Q(x,a)^{2})^{\frac{1}{2}}$.
@@
In Figure 1, we compare PID TD Learning with TD Learning when the gains are fixed and $\gamma=0.99$. In this case the acceleration depends on the choice of gains and the environment. We observe that we can achieve a drastic acceleration in Cliff Walk, and a minor acceleration in Chain Walk. We further investigate the speed-up achieved by PID TD Learning in Cliff Walk. In Figure 2, we observe that with Gain Adaptation and $\gamma=0.999$, we achieve a significant acceleration without the need to tune the controller gains. Figure 2 also shows how Gain Adaptation has modified the gains from their initial values.
@@
Figure 2: PID TD Learning with Gain Adaptation in Cliff Walk with $\gamma=0.999$. (Left) Comparison of value errors of PID TD Learning with TD Learning. Each curve is averaged over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 3: PID Q-Learning with Gain Adaptation in Chain Walk with $\gamma=0.999$. (Left) Comparison of value errors of PID Q-Learning with Q-Learning. Each curve is averaged over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 4: Comparison of PID Accelerated algorithms with the conventional ones for PE (Left) and Control (Right) problems in randomly generated Garnet environments with $\gamma=0.99$. Each curve is an average of 80 MDPs, run for 80 times each. Shaded area shows standard error.
@@ 7 Related Work
There is a growing literature of applying acceleration techniques to RL. Similar to PID VI, which as we showed leads to PID TD Learning and PID Q-Learning, many accelerated dynamic programming methods have closely related RL algorithms. The idea of momentum has been used for faster dynamic programming algorithms by Vieillard et al. (2020); Goyal and Grand-Clément (2023) and in RL setting has led to Speedy Q-Learning (Ghavamzadeh et al., 2011) and Momentum Q-Learning (Bowen et al., 2021). Zap Q-Learning (Devraj and Meyn, 2017) is another accelerated variant of Q-Learning based on second-order optimization methods. Anderson acceleration (Anderson, 1965) has been used for Anderson VI (Geist and Scherrer, 2018) and Anchoring acceleration (Halpern, 1967) is used in Anchord VI (Lee and Ryu, 2023). Matrix splitting is used to derive Operator Splitting VI (OSVI) (Rakhsha et al., 2022) and Deflated Dynamics VI (DDVI) (Lee et al., 2024), which are both extended to the RL setting through stochastic approximation. Recently, Rakhsha et al. (2024) has introduced the Model Correcting VI (MoCoVI) and Model Correcting Dyna (MoCoDyna) algorithms that achieve acceleration through model correction.
Gain adaptation in general has a long history in RL and closely-related literature. Kesten (1958) used an adaptive mechanism in the context of stochastic approximation in the 1950s. They describe a method for choosing the learning rate of SA that is very similar to the P component of the gain adaptation procedure we naturally derive. However, the algorithm is ad hoc in nature, and is not compatible with function approximation in any natural way. First order methods of adapting hyperparameters have been proposed, including IDBD (Sutton, 1992), the recent RL focused variant TIDBD (Kearney et al., 2018), and SMD (Schraudolph, 1999) which all tune learning rates by finding the gradient with respect to the history of errors. We refer to Sutton (2022) for a more in-depth history and overview of such techniques. These approaches are limited to controlling only the learning rate of the procedure and thus only attacking the error from sampling, not the bootstrapping error.
@@ 8 Conclusion
We showed how recent advances in accelerated planning and dynamic programming, specifically the PID Value Iteration algorithm, can be used to design algorithms for the RL setting. The proposed PID TD Learning and PID Q-Learning algorithms are accompanied by a gain adaptation mechanism, which tunes their hyperparameters on the fly. We provided theoretical analysis as well as empirical studies of these algorithms.
One limitation of the current work is that the proposed algorithms are only developed for finite MDPs where the value function, and all relevant quantities, can be represented exactly. For large MDPs, for example with continuous state spaces, we need to use function approximation. Developing PID TD Learning and PID Q-Learning with function approximation is therefore one important future direction. Another limitation of this work is that the gain adaptation procedure, even though empirically reliable, does not come with a convergence guarantee. Moreover, small changes in its hyperparameters, such as its meta-learning rate $\eta$, can cause large changes in the trajectory the value function takes during training. Another interesting research direction is then to develop a gain adaptation procedure that is less sensitive to the choice of hyperparameters and has a convergence guarantee. Finally, this work shows that the dynamics of RL can be significantly influenced by the PID controller, one of the simplest controllers in the arsenal of control engineering. Developing Planning and RL algorithms based on more sophisticated controllers is another promising research direction.
@@
In this section, we present the proof of Theorem 1. We first present some notation, and then the assumptions on the learning rate schedule $\mu$ and sequence of visited samples $(X_{t})_{t\geq 0}$. Let $r^{\pi}\in\mathbb{R}^{n}$ be the vector of the expected immediate rewards of following the policy at each state.
@@
Examples of learning rate schedules that satisfy Assumption 1 includes $\mu(t)=\frac{1}{t+1}$. The next assumption is on the balanced updates of states.
@@
We show the claim by applying the result by Borkar and Meyn (2000, Theorem 2.5) to our algorithm. We describe how PID TD Learning (3) is a special case of a convergent asynchronous stochastic approximation in (Borkar and Meyn, 2000). PID TD Learning updates three entries of $\tilde{V}_{t}$ at each iteration. Therefore, our set of indices that are updated (noted by $Y(n)$ in the original paper) is $Y_{t}\triangleq\{(X_{t},\texttt{v}),(X_{t},\texttt{z}),(X_{t},\texttt{v'})\}$’ ) }. The number of times the value for $(x,\texttt{f})\in{\mathcal{X}}\times\{\texttt{v},\texttt{z},\texttt{v'}\}$’ } is updated (noted by $\nu(i,n)$ in the original paper) is $N_{t}(x)$ and the communication delays are zero in our case. With these choices, the asynchronous stochastic approximation of (Borkar and Meyn, 2000, Equation 2.8) becomes
$\tilde{V}_{t+1}(X_{t},\texttt{f})\leftarrow\tilde{V}_{t}(X_{t},\texttt{f})+\mu(N_{t}(X_{t}))f[\tilde{V}_{t},D_{t}](X_{t},\texttt{f})\quad(\forall\texttt{f}\in\{\texttt{v},\texttt{z},\texttt{v'}\}),$’ } ) , (13)
and the other entries in $\tilde{V}_{t+1}$ remain the same as $\tilde{V}_{t}$. Here, $D_{t}\in{\mathcal{D}}$ for all $t$ are independently and identically distributed (i.i.d.), and $f\colon\mathbb{R}^{3n}\times{\mathcal{D}}\to\mathbb{R}^{3n}$ can be an arbitrary function. To obtain PID TD Learning in this form, we define ${\mathcal{D}}\triangleq({\mathcal{X}}\times[0,1])^{n}$. For any $t$, we choose $D_{t}$ to be a dataset $\{(R_{x,t},X^{\prime}_{x,t})\}_{x\in{\mathcal{X}}}$ where for each state $x$ contains the random reward $R_{x,t}\sim\mathcal{R}^{\pi}(x)$ and $X^{\prime}_{x,t}\sim{\mathcal{P}}^{\pi}(\cdot|x)$. Then we define $f$ such that for all $x\in{\mathcal{X}}$,
$f[\tilde{V}_{t},D_{t}](x,\texttt{v})\triangleq\kappa_{p}b_{x,t}+\kappa_{I}(\beta z_{t}(x)+\alpha b_{x,t})+\kappa_{d}(V_{t}(x)-V^{\prime}_{t}(x)),$
$f[\tilde{V}_{t},D_{t}](x,\texttt{z})\triangleq\beta z_{t}(x)+\alpha b_{x,t}-z_{t}(x),$
$f[\tilde{V}_{t},D_{t}](x,\texttt{v'})\triangleq V_{t}(x)-V^{\prime}_{t}(x).$
where $b_{x,t}=R_{x,t}+\gamma V_{t}(X^{\prime}_{x,t})-V_{t}(x)$. This yields the exact same PID TD Learning updates.
The function $h(\tilde{V})=\mathbb{E}[f(\tilde{V},D_{1})]$ in (Borkar and Meyn, 2000) is equal to $L_{g}^{\pi}\tilde{V}-\tilde{V}$ in our setting. Note that $h$ is Lipschitz since it is an affine linear operator. The function $h_{\infty}(\tilde{V})$ exists and is equal to $(A_{g}^{\pi}-I)\tilde{V}$. Therefore, we require the origin point to be an asymptotically stable equilibrium of the ODE
$\dot{u}(t)=h_{\infty}(u(t))=(A_{g}^{\pi}-I)u(t),$
which is satisfied due to the assumption on the eigenvalues of $A_{g}^{\pi}$ and the fact that the solution of the above ODE is $\exp{[(A_{g}^{\pi}-I)t]}u_{0}$ (Teschl, 2012) for any starting point $u_{0}$.
@@
$\dot{u}(t)=h(u(t))=L_{g}^{\pi}u(t)-u(t)=(A_{g}^{\pi}-I)u(t)+b_{g}^{\pi}.$
is $-(A_{g}^{\pi}-I)^{-1}b_{g}^{\pi}$ which is equal to $\tilde{V}^{\pi}$ due to $\tilde{V}^{\pi}=L_{g}^{\pi}\tilde{V}^{\pi}=A_{g}^{\pi}\tilde{V}^{\pi}+b_{g}^{\pi}$.
Finally, note that Lemma 1 established the required property of the noise. The remaining assumptions of Borkar and Meyn (2000) are satisfied due to Assumption 1 and 2.
@@
The existence of $S^{\prime}$ such that $\left\lVert A\right\rVert_{2,S^{\prime}}\leq\rho(A)+\delta$ is a consequence of the proof of Theorem 4.4 in Householder (1958). Due to equivalence of norms in finite dimensions, there exists $u>0$ such that for any $v\in\mathbb{R}^{d}$, $\left\lVert v\right\rVert_{2,S^{\prime}}\leq u\left\lVert v\right\rVert_{\infty}$. Define $S=\frac{1}{u}S^{\prime}$. Consequently, $\left\lVert v\right\rVert_{2,S}\leq\left\lVert v\right\rVert_{\infty}$ for any $v$. Now from Theorem 2.10 in Householder (1958), we have $\left\lVert A\right\rVert_{2,S^{\prime}}=\left\lVert S^{\prime}AS^{\prime-1}\right\rVert_{2}$ and $\left\lVert A\right\rVert_{2,S}=\left\lVert SAS^{-1}\right\rVert_{2}$ which means
@@
Assume a dataset $\{(x,R_{x},X^{\prime}_{x})\}_{x}$ is given, where for each state $x$ contains the random reward $R_{x}\sim\mathcal{R}^{\pi}(x)$ and $X^{\prime}_{x}\sim{\mathcal{P}}^{\pi}(\cdot|x)$, and $\pi$ is $d$-deterministic in the environment. For an arbitrary value function $V$, define $W\colon{\mathcal{X}}\to\mathbb{R}$ as
$W(x)\triangleq R_{x}+\gamma V(X^{\prime}_{x})-r^{\pi}(x)-\gamma({\mathcal{P}}^{\pi}V)(x).$
@@
The result follows from the fact that for any random vector $Z=[Z_{1},\ldots,Z_{k}]^{\top}$,
$\mathbb{E}\left[\left\lVert Z\right\rVert_{\infty}^{2}\right]\leq\mathbb{E}\left[\sum_{i}Z_{i}^{2}\right]=\sum_{i}E\left[Z_{i}^{2}\right].$